# Memory in neural activity: long-range order without criticality

### v4

@@ I Introduction
Fig. 1: An illustration of our phenomenological model. The bottom image depicts our 2D square lattice network, meant to emulate the cerebral cortex, where neural activities $\rho_{i}$ are represented by blue cubes. Each cube has some amount of available resources $R_{i}$, represented by the red filling and acting as memory degrees of freedom. The green lines represent diffusive couplings between any activity region and its four nearest-neighbors. The top image zooms in on one region; the orange arrows show the dynamical coupling between $\rho_{i}$ and $R_{i}$, and the flow fields $F_{\rho_{i}}$ and $F_{R_{i}}$ can be read from the right hand side of Eqs. 1 and 2, respectively.
@@ II Model
Our mesoscopic cortical dynamics model is illustrated heuristically in Fig. 1. At each site in a two-dimensional (2D) square lattice network of size $L^{2}$ exists one region of neural activity, the smallest functional unit in our model. These regions are constructed by applying a Landau-Ginzburg approach [22, 23] to the widely-accepted Wilson-Cowan model for neural activity [20]. Individual regions of activity $\rho_{i}$, each consisting of thousands of individual neurons, are taken to follow a Wilson-Cowan equation. We take the region-dependent resource availability $R_{i}$ dynamics to be similar to those of the continuous time Tsodyks-Markram model for synaptic plasticity [24]. The activity regions are coupled diffusively, and stochasticity is introduced in the form of additive noise, both in line with the Landau-Ginzburg approach:
$\dot{\rho_{i}}(t)$ $=(-a+R_{i})\rho_{i}+b\rho_{i}^{2}-c\rho_{i}^{3}+h+D\nabla^{2}\rho_{i}+\sigma\eta_{i}\,$$\quad\;\;$ (1)
$\dot{R_{i}}(t)$ $=\delta-\frac{1}{\tau_{D}}(R_{i}\rho_{i}+\sigma\eta_{i})\,$$\quad\;\;$ (2)
@@
where $i=1,2,\dots,L^{2}$. In Eqs. 1 and 2, $\eta_{i}$ corresponds to zero-mean, unit-variance Gaussian white noise.
[…]
@@
We emphasize that Eqs. 1 and 2 are very similar to the ones employed by di Santo et al. [21] in their study of the emergence of scale-free avalanches at the edge of synchronization. However, their study does not discuss or identify the role of memory in the generation of an LRO phase. In addition, to more directly focus on the role memory plays in inducing LRO, our dynamics feature a few changes from the model in [21]. First, our noise term $\sigma\eta_{i}$ is additive, rather than multiplicative. Second, we introduce an identical source of Gaussian white noise in the resource dynamics, to better reflect the stochasticity of real physical systems. Third, we represent the replenishing of resources by a constant parameter $\delta=0.004$, rather than a function of $R_{i}$. And fourth, we have decided to scale the noise term in the resource dynamics by the value of $\tau_{D}$, which ensures the noise does not dominate the resource dynamics when the product $R_{i}\rho_{i}$ is small and thus that $\tau_{D}$ is representative of a resource decay timescale for arbitrary noise strength $\sigma$. Despite these changes, we obtain phase structures similar to those reported in [21] when other parameters are kept at the same values, giving us confidence that these modifications do not significantly influence our conclusions. In fact, as we will show in Sec. III, the nearest-neighbor diffusive couplings, $D\nabla^{2}\rho_{i}$, and the activity-resource couplings, $R_{i}\rho_{i}$, are the dominant factors in inducing LRO. Our numerical results in IV also support this conclusion.
III Analytical results
@@
$\left\{\begin{aligned} \dot{\rho}_{i}(t)\rightarrow\dot{\rho}_{i}(t)&=R_{i}\rho_{i}+D\nabla^{2}\rho_{i}\;,\quad\quad\quad\quad\\ \dot{R}_{i}(t)\rightarrow\dot{R}_{i}(t)&=-\frac{1}{\tau_{D}}R_{i}\rho_{i}\;,\quad\quad\quad\quad\end{aligned}\right$. (3)
@@
To this end, we will perform a time coarse-graining approach inspired by [2] to evaluate $\rho_{i}(t)$ and $R_{i}(t)$. Time is doubly discretized: first, coarsely over the timescale of the activity (memory) dynamics, $\tau_{D}$, and then more finely over the neural activity timescale $1/D$. Fig. 2 clarifies the discretization procedure. First, we evolve $\rho_{i}(t)$ over one resource timescale $\tau_{D}$, so $R_{i}(t)$ can be treated as approximately constant:
$\rho_{i}(\tau_{D})$ $=\rho_{i}(0)+\int_{0}^{\tau_{D}}dt\dot{\rho}_{i}(t)$ (4)
$\approx\rho_{i}(0)+\tau_{D}\big{(}R_{i}(0)\overline{\rho_{i,0}}+D\nabla^{2}\overline{\rho_{i,0}}\big{)}.$
Using similar notation as in [2], we define $\overline{\rho_{i,k}}$ as
$\overline{\rho_{i,k}}\equiv\frac{1}{n}\sum_{p=0}^{n-1}\rho_{i}(k\tau_{D}+p/D),$ (5)
[…]
If we evaluate $\rho_{i}(t)$ at some time $t>\tau_{D}$, the resources have had time to update, so their dynamics must now be considered. For example, when $t=2\tau_{D}$, we will first need to find $R_{i}(\tau_{D})$ using a similar method:
@@
$\rho_{i}(2\tau_{D})$ $\approx\rho_{i}(\tau_{D})+\tau_{D}\big{(}R_{i}(\tau_{D})\overline{\rho_{i,1}}+D\nabla^{2}\overline{\rho_{i,1}}\big{)}$ (7)
$\approx\rho_{i}(\tau_{D})+\tau_{D}\big{(}R_{i}(0)(1-\overline{\rho_{i,0}})\overline{\rho_{i,1}}+D\nabla^{2}\overline{\rho_{i,1}}\big{)}.$
[…]
Typically, this diffusive effect alone would decay exponentially and would not suggest LRO. However, because $\overline{\rho_{i,0}}$ feeds back into the dynamics for $\rho_{i}(2\tau_{D})$ explicitly, beyond-nearest-neighbor couplings between activities in the lattice are created. These couplings are second order, but it can be shown that $N^{\text{th}}$ order couplings will be induced after time $N\tau_{D}$ has passed. We provide these generalizations along with more detailed justifications in Sec. A of the Supplementary Information (SI). Therefore, after a sufficiently long time, we anticipate that regions of activity any arbitrary distance from one another in the lattice will become coupled.
@@ IV Numerical results
Fig. 3: Illustration of the dynamics of neural activity $\rho$ and avalanche size distributions for different values of memory time scale $\tau_{D}$. (a) Raster plots of $\rho_{i}(t)$ in a $N=64^{2}$ lattice of activities. Insets depict 2D snapshots of the dynamics at a specific time. Videos of the temporal dynamics of this network for various $\tau_{D}$ can be found in the SI. As the value of $\tau_{D}$ increases, the system undergoes phase changes from the down phase (leftmost panel), to the LRO phase (middle 3 panels), to the rigid phase (rightmost panel). (b) Avalanche size distribution in the down phase ($\tau_{D}=15$). (c-e) Avalanche size distributions in the LRO phase ($\tau_{D}=25,51,77$). The insets are finite-size scaling curves transformed according to Eq. 9, showing that the biases from ideal power-law distributions stem from finite-size effects. The additional peaks in the distributions in (d) and (e) correspond to avalanches that propagate back through the lattice more than once. (f) Avalanche size distribution in the rigid (up) phase ($\tau_{D}=88$). For $L=\{64,128,256,512\}$, the statistics from each ensemble are taken from $\{100,400,2000,4800\}$ instances, with each instance simulated for $\{50,20,5,1.25\}\times 10^{4}$ time steps, respectively. Parameter values: $a=1$, $b=1.5$, $c=1$, $h=10^{-7}$, $D=1$, $\sigma=0.1$, $\delta=0.004$, $\epsilon=0.5$, $\Delta t=0.01$, and $\Delta_{tw}=0.3$.
To show explicitly the anticipated emergence of LRO in our cortical dynamics model, we simulate Eqs. 1 and 2 while varying the resource decay timescale, $\tau_{D}$. Fig. 3(a) depicts raster plots that visualize the neural activity $\rho(t)$ across the whole lattice (for N = $64^{2}$) for 5 different values of $\tau_{D}$, with the insets featuring 2D snapshots of the dynamics. Higher activity levels are indicated by darker blues. Each horizontal slice in the raster plot represents the activity level of a single region $\rho_{i}(t)$. By varying $\tau_{D}$, with all other parameters fixed, three distinct signal patterns of $\rho(t)$ can be observed: a “down” phase, an intermediate phase, and an “up” phase.
@@ IV.1 Avalanche Distributions
To further characterize the neural activity dynamics in each phase, we extract avalanche distributions from the neural activity levels $\rho_{i}(t)$. See Sec. VI.2 for more details on how we define and calculate avalanche size/duration distributions. We can then study how the distributions of avalanche sizes/durations change as $\tau_{D}$ is changed to further characterize each phase as a function of the resource depletion timescale. Avalanche size distributions are plotted in Fig. 3(b-f). Since the avalanche duration distributions look heuristically very similar to the avalanche size distributions, we choose to include the latter in Fig. 3 of Sec. B.3 of the SI. All following discussions of avalanche size/duration distributions can be applied interchangeably. All distributions are represented as histograms whose bin width is determined using Scott’s normal reference rule [25].
@@
Above, $\alpha_{s}$ and $\alpha_{T}$ are the exponents for the avalanche size and duration distributions, respectively. $\gamma$ is defined via $\langle s\rangle(T)=T^{\gamma}$, in which, notably, the mean size of avalanches as a function of avalanche duration also follows a power law.
@@ IV.2 Finite-Size Scaling
where $v$ = s or T depending on the type of distribution we consider (spatial or temporal, respectively), N = $L^{2}$ is the system size, and parameters $\alpha$ and $\beta$ are the critical and cutoff exponents, respectively.
@@ IV.3 Width of the LRO Phase and Correlation Length
Fig. 4: The correlation length, $\xi$, defined as the average distance between two units within the same avalanche (excluding system-wide avalanches) [28], as a function of $\tau_{D}$. For a large range of $\tau_{D}$ ($25\lesssim\tau_{D}\lesssim 82$), the correlation length extends far beyond nearest-neighbors, indicating the existence of a robust phase of LRO.
Having calculated the avalanche size distributions, we define the correlation length, $\xi$, as the average distance between two sites within the same avalanche, excluding system-wide avalanches. This definition follows the convention in percolation theory [28], and a detailed equation for computing $\xi$ can be found in Sec. VI.1.
When plotting the correlation length $\xi$ against $\tau_{D}$, as shown in Fig. 4, we observe that $\xi(\tau_{D})>>1$ (in units of lattice spacing) within the range $25\leq\tau_{D}\leq 82$. This implies that the correlation length extends far beyond nearest-neighbors. Given that the avalanches also exhibit consistent power-law distributions, we conclude that in the thermodynamic limit, $\xi$ effectively diverges within much of the central region, further indicating that a robust phase of LRO exists.

### v5

@@ I Introduction
Fig. 1: An illustration of our phenomenological model. The bottom image depicts our 2D square lattice network, meant to emulate the cerebral cortex, where neural activities $\rho_{\vec{x}}$ are represented by blue cubes. Each cube has some amount of available resources $R_{\vec{x}}$, represented by the red filling and acting as memory degrees of freedom. The green lines represent diffusive couplings between any activity region and its four nearest-neighbors. The top image zooms in on one region; the orange arrows show the dynamical coupling between $\rho_{\vec{x}}$ and $R_{\vec{x}}$, and the flow fields $F_{\rho}$ and $F_{R}$ can be read from the right hand side of Eqs. 1 and 2, respectively.
@@ II Model
Our mesoscopic cortical dynamics model is illustrated heuristically in Fig. 1. At each site in a two-dimensional (2D) square lattice network of size $L^{2}$ exists one region of neural activity, the smallest functional unit in our model. These regions are constructed by applying a Landau-Ginzburg approach [22, 23] to the widely-accepted Wilson-Cowan model for neural activity [20]. Individual regions of activity $\rho_{\vec{x}}$, each consisting of thousands of individual neurons, are taken to follow a Wilson-Cowan equation. We take the region-dependent resource availability $R_{\vec{x}}$ dynamics to be similar to those of the continuous time Tsodyks-Markram model for synaptic plasticity [24]. The activity regions are coupled diffusively, and stochasticity is introduced in the form of additive noise, both in line with the Landau-Ginzburg approach:
$\dot{\rho}_{\vec{x}}(t)$ $=(-a+R_{\vec{x}})\rho_{\vec{x}}+b\rho_{\vec{x}}^{2}-c\rho_{\vec{x}}^{3}+h$ (1)
$\hskip 99.58464pt+D\nabla^{2}\rho_{\vec{x}}+\sigma\eta_{\vec{x}}\,$$\quad\;\;$
$\dot{R}_{\vec{x}}(t)$ $=\delta-\frac{1}{\tau_{D}}(R_{\vec{x}}\rho_{\vec{x}}+\sigma\zeta_{\vec{x}})\,$$\quad\;\;$ (2)
@@
where ${\vec{x}}=(x_{1},x_{2})$ and $x_{i}=1,2,\dots,L$. In Eqs. 1 and 2, $\eta_{\vec{x}}$ and $\zeta_{\vec{x}}$ both correspond to zero-mean, unit-variance Gaussian white noise.
The neural activity Eq. 1 aims to emulate the time evolution of the neural voltage. Diffusivity exists here in the form of $D\nabla^{2}\rho_{\vec{x}}=D\sum_{\vec{y}\in n.n.\vec{x}}(\rho_{\vec{y}}-\rho_{\vec{x}})$, where $n.n.\vec{x}$ represents the nearest-neighbors to a lattice position $\vec{x}$. Note that the lattice spacing has been normalized to $1$, so $D$ is a diffusive frequency. Following [21], we have chosen the following values for the parameters: $D=1$, $a=1$, $b=1.5$, $c=1$, $h=10^{-7}$, and $\sigma=0.1$. A detailed explanation of the remaining terms (in particular, the effects of coefficients $a$, $b$, $c$, and $h$) can be found in [21]. The resource availability Eq. 2 holistically represents the available resources (e.g., ions, neurotransmitters, ATP, etc.) that may be consumed by the activity region at site $\vec{x}$. For our study, we consider the time scale $\tau_{D}$ as a control parameter, as it governs the rate of resource decay with respect to the fixed diffusive timescale $\sim 1/D$.
@@
We emphasize that Eqs. 1 and 2 are very similar to the ones employed by di Santo et al. [21] in their study of the emergence of scale-free avalanches at the edge of synchronization. However, their study does not discuss or identify the role of memory in the generation of an LRO phase. In addition, to more directly focus on the role memory plays in inducing LRO, our dynamics feature a few changes from the model in [21]. First, we introduce an identical source of Gaussian white noise in the resource dynamics, to better reflect the stochasticity of real physical systems. Second, our noise terms $\sigma\eta_{\vec{x}}$ and $\sigma\zeta_{\vec{x}}$ are additive, rather than multiplicative. Third, we represent the replenishing of resources by a constant parameter $\delta=0.004$, rather than a function of $R_{\vec{x}}$. And fourth, we have decided to scale the noise term in the resource dynamics by the value of $\tau_{D}$, which ensures the noise does not dominate the resource dynamics when the product $R_{\vec{x}}\rho_{\vec{x}}$ is small and thus that $\tau_{D}$ is representative of a resource decay timescale for arbitrary noise strength $\sigma$. Despite these changes, we obtain phase structures similar to those reported in [21] when other parameters are kept at the same values, giving us confidence that these modifications do not significantly influence our conclusions. In fact, as we will show in Sec. III, the nearest-neighbor diffusive couplings, $D\nabla^{2}\rho_{\vec{x}}$, and the activity-resource couplings, $R_{\vec{x}}\rho_{\vec{x}}$, are the dominant factors in inducing LRO. Our numerical results in IV also support this conclusion.
III Analytical Understanding
@@
$\left\{\begin{aligned} \dot{\rho}_{\vec{x}}(t)\rightarrow\dot{\rho}_{\vec{x}}(t)&=R_{\vec{x}}\rho_{\vec{x}}+D\nabla^{2}\rho_{\vec{x}}\;,\quad\quad\quad\quad\\ \dot{R}_{\vec{x}}(t)\rightarrow\dot{R}_{\vec{x}}(t)&=-\frac{1}{\tau_{D}}R_{\vec{x}}\rho_{\vec{x}}\;,\quad\quad\quad\quad\end{aligned}\right$. (3)
@@
To this end, we will perform a time coarse-graining approach inspired by [2] to evaluate $\rho_{\vec{x}}(t)$ and $R_{\vec{x}}(t)$. Time is doubly discretized: first, coarsely over the timescale of the activity (memory) dynamics, $\tau_{D}$, and then more finely over the neural activity timescale $1/D$. Fig. 2 clarifies the discretization procedure. First, we evolve $\rho_{\vec{x}}(t)$ over one resource timescale $\tau_{D}$, so $R_{\vec{x}}(t)$ can be treated as approximately constant:
$\rho_{\vec{x}}(\tau_{D})$ $=\rho_{\vec{x}}(0)+\int_{0}^{\tau_{D}}dt\dot{\rho}_{\vec{x}}(t)$ (4)
$\approx\rho_{\vec{x}}(0)+\tau_{D}\big(R_{\vec{x}}(0)\overline{\rho_{\vec{x},0}}+D\nabla^{2}\overline{\rho_{\vec{x},0}}\big).$
Using similar notation as in [2], we define $\overline{\rho_{\vec{x},k}}$ as
$\overline{\rho_{\vec{x},k}}\equiv\frac{1}{n}\sum_{p=0}^{n-1}\rho_{\vec{x}}(k\tau_{D}+p/D),$ (5)
which is just the discretized average value of the activity at site $\vec{x}$ over the interval $k\tau_{D},(k+1)\tau_{D})$. Here, $n\equiv\lfloor\tau_{D}D\rfloor$ is an approximate ratio between characteristic timescales (or, equivalently, the number of activity timesteps in a single resource timestep, with $\lfloor\cdot\rfloor$ the floor function). This averaging is justified in a similar manner to the $R_{\vec{x}}$ averaging: since the activity variables do not change much over timescales less than $1/D$, we can treat them as approximately constant during intervals of length $1/D$.
If we evaluate $\rho_{\vec{x}}(t)$ at some time $t>\tau_{D}$, the resources have had time to update, so their dynamics must now be considered. For example, when $t=2\tau_{D}$, we will first need to find $R_{\vec{x}}(\tau_{D})$ using a similar method:
@@
$\rho_{\vec{x}}(2\tau_{D})$ $\approx\rho_{\vec{x}}(\tau_{D})+\tau_{D}\big(R_{\vec{x}}(\tau_{D})\overline{\rho_{\vec{x},1}}+D\nabla^{2}\overline{\rho_{\vec{x},1}}\big)$ (7)
$\approx\rho_{\vec{x}}(\tau_{D})+\tau_{D}\big(R_{\vec{x}}(0)(1-\overline{\rho_{\vec{x},0}})\overline{\rho_{\vec{x},1}}+D\nabla^{2}\overline{\rho_{\vec{x},1}}\big).$
Explicitly, we see that activity-activity couplings have been induced by the presence of memory in the term $\tau_{D}R_{\vec{x}}(0)\overline{\rho_{\vec{x},0}}\,\overline{\rho_{\vec{x},1}}$. While this correlation appears to be only temporal, $\overline{\rho_{\vec{x},0}}$ and $\overline{\rho_{\vec{x},1}}$ implicitly depend on many of the sites around the lattice position $\vec{x}$ via diffusion. Since $\tau_{D}\gtrsim 1/D$, the diffusion will cause any $\rho_{\vec{x}}(\tau_{D})$ to depend on initial activity values up to $n=\lfloor\tau_{D}D\rfloor$ lattice sites away.
Typically, this diffusive effect alone would decay exponentially and would not suggest LRO. However, because $\overline{\rho_{\vec{x},0}}$ feeds back into the dynamics for $\rho_{\vec{x}}(2\tau_{D})$ explicitly, beyond-nearest-neighbor couplings between activities in the lattice are created. These couplings are second order, but it can be shown that $N^{\text{th}}$ order couplings will be induced after time $N\tau_{D}$ has passed. We provide these generalizations along with more detailed justifications in Sec. A of the Supplementary Information (SI). Therefore, after a sufficiently long time, we anticipate that regions of activity any arbitrary distance from one another in the lattice will become coupled.
@@ IV Numerical Results
Fig. 3: Illustration of the dynamics of neural activity $\rho$ and avalanche size distributions for different values of memory time scale $\tau_{D}$. (a) Raster plots of $\rho_{\vec{x}}(t)$ in a $N=64^{2}$ lattice of activities. Each horizontal line represents the temporal evolution of one site, while the insets show corresponding 2D snapshots of spatial activity patterns at representative times. Videos of the temporal dynamics of this network for various $\tau_{D}$ can be found in the SI. As the value of $\tau_{D}$ increases, the system undergoes phase changes from the down phase (leftmost panel), to the LRO phase (middle 3 panels), to the rigid phase (rightmost panel). (b) Avalanche size distribution in the down phase ($\tau_{D}=15$). (c-e) Avalanche size distributions in the LRO phase ($\tau_{D}=25,51,77$). The insets are finite-size scaling curves transformed according to Eq. 9, showing that the biases from ideal power-law distributions stem from finite-size effects. The additional peaks in the distributions in (d) and (e) correspond to avalanches that propagate back through the lattice more than once. (f) Avalanche size distribution in the rigid (up) phase ($\tau_{D}=88$). For $L=\{64,128,256,512\}$, the statistics from each ensemble are taken from $\{100,400,2000,4800\}$ instances, with each instance simulated for $\{50,20,5,1.25\}\times 10^{4}$ time steps, respectively. Parameter values: $a=1$, $b=1.5$, $c=1$, $h=10^{-7}$, $D=1$, $\sigma=0.1$, $\delta=0.004$, $\epsilon=0.5$, $\Delta t=0.01$, and $\Delta_{tw}=0.3$.
To show explicitly the anticipated emergence of LRO in our cortical dynamics model, we simulate Eqs. 1 and 2 while varying the resource decay timescale, $\tau_{D}$. Fig. 3(a) depicts raster plots that visualize the neural activity $\rho(t)$ across the whole lattice (for N = $64^{2}$) for 5 different values of $\tau_{D}$, with the insets featuring 2D snapshots of the dynamics. Higher activity levels are indicated by darker blues. Each horizontal slice in the raster plot represents the activity level of a single region $\rho_{\vec{x}}(t)$. By varying $\tau_{D}$, with all other parameters fixed, three distinct signal patterns of $\rho(t)$ can be observed: a “down” phase, an intermediate phase, and an “up” phase.
@@ IV.1 Avalanche Distributions
To further characterize the neural activity dynamics in each phase, we extract avalanche distributions from the neural activity levels $\rho_{\vec{x}}(t)$. See Sec. VI.2 for more details on how we define and calculate avalanche size/duration distributions. We can then study how the distributions of avalanche sizes/durations change as $\tau_{D}$ is changed to further characterize each phase as a function of the resource depletion timescale. Avalanche size distributions are plotted in Fig. 3(b-f). Since the avalanche duration distributions look heuristically very similar to the avalanche size distributions, we choose to include the latter in Fig. 3 of Sec. B.3 of the SI. All following discussions of avalanche size/duration distributions can be applied interchangeably. All distributions are represented as histograms whose bin width is determined using Scott’s normal reference rule [25].
@@
Above, $\alpha_{s}$ and $\alpha_{T}$ are the critical exponents for the avalanche size and duration distributions, respectively. $\gamma$ is defined via $\langle s\rangle(T)=T^{\gamma}$, in which, notably, the mean size of avalanches as a function of avalanche duration $T$ also follows a power law.
@@ IV.2 Finite-Size Scaling
where $v$ = $s$ or $T$ depending on the type of distribution we consider (spatial or temporal, respectively), N = $L^{2}$ is the system size, and parameters $\alpha$ and $\beta$ are the critical and cutoff exponents, respectively. The critical exponent $\alpha$ is the effective power-law exponent the distribution follows, whereas the cutoff exponent scales the distribution according to finite-size effects.
@@ IV.3 Width of the LRO Phase and Correlation Length
Fig. 4: The correlation length, $\xi$, defined as the average distance between two units within the same avalanche (excluding system-wide avalanches), averaged over the entire distribution of avalanches [28], as a function of $\tau_{D}$. For a large range of $\tau_{D}$ ($25\lesssim\tau_{D}\lesssim 82$), the correlation length extends far beyond nearest-neighbors, indicating the existence of a robust phase of LRO.
Having computed the avalanche size distributions, we next define the correlation length, $\xi$, following the convention in percolation theory [28]. Specifically, we calculate the average distance between two units belonging to the same avalanche (excluding system-wide avalanches) and then average this quantity over the entire avalanche ensemble. The detailed formula for computing $\xi$ is provided in Sec. VI.1. This definition characterizes the typical spatial extent of an avalanche, which is inherently bounded by the system size $L$.
Fig. 4 shows $\xi$ as a function of $\tau_{D}$. Within the range $25\leq\tau_{D}\leq 82$, $\xi(\tau_{D})\gg 1$ (in units of lattice spacing), indicating correlations that extend well beyond nearest neighbors. In this regime, avalanches span a broad range of sizes, including events that approach the full lattice size, consistent with the emergence of long-range order (LRO). Moreover, $\xi$ increases systematically with $L$ in this phase, as larger lattices allow larger avalanches and therefore longer correlation lengths. This scaling behavior reflects the scale-free nature of the avalanche distributions: as $L$ increases, the upper cutoff of the power law extends, and in the thermodynamic limit ($L\to\infty$), $\xi$ is expected to diverge. This divergence signifies that the LRO phase is a genuine extended phase—robust across a wide parameter range—rather than a fine-tuned critical point.